Outose Concept manufactures small tables in its Processing Department. Direct materials are added at the initiation of the production cycle and must be bundled in single kits for each unit. Conversion costs are incurred evenly throughout the production cycle. Before inspection, some units are spoiled due to undetectable materials defects. Spoiled units generally constitute 3% of the good units. Data for December 2017 are as follows:

WIP, beginning inventory 12/1/2017                              22,700 units

        Direct materials (100% complete)

        Conversion costs (80% complete)

Started during December                                                  76,300 units

Completed and transferred out 12/31/2017                 72,100 units

WIP, ending inventory 12/31/2017                                 18,100 units

        Direct materials (100% complete)

        Conversion costs (60% complete)

Costs for December:

        WIP, beginning Inventory:

               Direct materials                                                          $151,000

               Conversion costs                                                             76,900

        Direct materials added                                                       231,400

        Conversion costs added                                                     294,000

What is the total cost per equivalent unit using the weighted-average method of process costing? (Round any cost per unit calculations to the nearest cent.)

A.

$3.69

B.

$4.04

C.

$7.90

D.

$3.86
Answer #1

Equivalent unit of material = 80900+18100 = 99000

Equivalent unit of conversion =80900+(18100*60%) = 91760

Cost per equivalent unit of material = (151000+231400)/99000 = 3.86

Cost per equivalent unit of conversion = (76900+294000)/91760 = 4.04

Total cost per equivalent unit = 4.04+3.86 = $7.90

So answer is c) $7.90